# ReviewFlow: Intelligent Scaffolding to Support Academic Peer Reviewing

### 05 Feb 2024 v1

@@ Abstract.
Peer review is a cornerstone of science. Research communities conduct peer reviews to assess contributions and to improve the overall quality of science work. Every year, new community members are recruited as peer reviewers for the first time. How could technology help novices adhere to their community’s practices and standards for peer reviewing? To better understand peer review practices and challenges, we conducted a formative study with 10 novices and 10 experts. We found that many experts adopt a workflow of annotating, note-taking, and synthesizing notes into well-justified reviews that align with community standards.Novices lack timely guidance on how to read and assess submissions and how to structure paper reviews. To support the peer review process, we developed ReviewFlow – an AI-driven workflow that scaffolds novices with contextual reflections to critique and annotate submissions, in-situ knowledge support to assess novelty, and notes-to-outline synthesis to help align peer reviews with community expectations. In a within-subjects experiment, 16 inexperienced reviewers wrote reviews using ReviewFlow and a baseline environment with minimal guidance. Participants produced more comprehensive reviews using ReviewFlow than the baseline, calling out more pros and cons, but they still struggled to provide actionable suggestions to address the weaknesses. While participants appreciated the streamlined process support from ReviewFlow, they also expressed concerns about using AI as part of the scientific review process. We discuss the implications of using AI to scaffold peer review process on scientific work and beyond.
@@
copyright: acmcopyrightjournalyear: 2024price: 15.00ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Peer review is a cornerstone of academic research, ensuring the quality, credibility, and reliability of scientific research (Price and Flach, 2017). The peer review process seeks to assess whether submissions contribute new knowledge to a research community and generate feedback that helps authors improve the quality of their work  (Jefferson et al., 2002; Shah, 2022). Many communities are seeing a rapid increase in submissions (McCook, 2006; Stelmakh et al., 2019; Shah, 2022; Arous et al., 2021; Stelmakh et al., 2021a); while this could be seen as an indicator of scientific progress, it also has increased the pressure on peer review ecosystems. To meet increased demand, many research communities recruit a significant number of first-time reviewers for each review cycle. For example, in 2023, the ACM CHI Conference on Human Factors in Computing Systems reported that more than 50% of ACs who reviewed papers are first-time ACs (chi, 2015).
@@
Peer reviewing is a complicated and challenging task. Reviewers need to understand the paper, evaluate the scientific content using domain knowledge, make a fair decision, and compose a comprehensive draft to communicate their assessments and recommendations (Shah, 2022). It is a task that requires a deep understanding of the subject, the critical thinking process, and the ability to provide constructive feedback. Conferences and journals need to ensure that first-time reviewers meet the standards of the research community.
@@
Figure 1. ReviewFlow interface. Users can (A) receive section level reflection cues guided by community criteria; (B) request phase level reflection cues adapted to the highlight paper content; (C) click the citation to check in-situ summarization; (D) check the recommended citations that are potentially missed by the paper; (E) click to summarize the notes into high level structured outline and expand it to a detailed outline
@@
One approach to preparing novices for any complex task is to provide scaffolding, a strategy used in educational settings to support learning and mastery (Reiser, 2004; Collins, 2006). Scaffolding takes many forms, including examples of prior work, templates, or hints to help novices think more and potentially perform more on par with experts (Holton and Clarke, 2006; Saye and Brush, 2002). Previous studies found that scaffolded examples and templates can even help learners perform similarly to experts in terms of feedback quality (Yuan et al., 2016). Recently, LetterSmith’s instructional approach to aid writing called “scaffolded annotation” which provided key components of the writing tasks and annotated expert examples, helped professional writing students improve the quality of their early-stage drafting in a new genre (Hui and Sprouse, 2023). Advances in AI and large-language models (LLMs) in particular, have the potential to make scaffolding even more effective, because they can interpret and adapt to the user’s work context (Rachatasumrit et al., 2022). For example, the CReBot project leverages LLMs to generate guiding questions to help induce creative thinking while reading academic papers (Peng et al., 2022). Experimental results showed that CReBot can engage novices in reading and help them comprehend the paper content. While LLMs show great potential, they are also known to hallucinate (Bang et al., 2023) and perpetuate systemic biases (Huang et al., 2019; Jakesch et al., 2023) that could negatively impact the effectiveness of scaffolding. While intelligent scaffolding has proven valuable for well-scoped tasks, academic peer review involves reading submissions with a critical eye, synthesizing knowledge, and making a well-justified judgment for acceptance or rejection. Our research investigates how intelligent scaffolds can guide a complex, multi-faceted workflow for academic peer reviewing without biasing the ultimate decisions. This paper explores two key research questions:
  • What are the strategies adopted by experts and challenges faced by novices during the peer review process?
  • How can we leverage LLMs to intelligently scaffold the peer review process for novices without introducing bias or over-reliance?
@@
Based on these insights, we developed a prototype – “ReviewFlow” [^1] ReviewFlow Code Repository:https://github.com/creativecolab/ReviewFlow.git, a platform for writing peer reviews that incorporate intelligent scaffolding to support a workflow for inexperienced reviewers. ReviewFlow incorporates a range of features to facilitate the review process: (1) Contextual cues are embedded questions that help reviewers reflect on the paper. ReviewFlow includes section-level cues guided by the community’s review criteria as well as phrase-level cues that adapt to the paper’s content. (2) In-situ citation recommendations show relevant but non-cited papers that may help reviews assess the novelty compared to existing work. (3) Notes-to-outline synthesis guides reviewers to organize notes and structure reviews to align with community standards. ReviewFlow gathers all the notes left by the reviewer and leverages an LLM to summarize notes into a high-level outline. Reviewers can revise and add detail to the outline while adhering to community standards.
We conducted a within-subjects study to evaluate ReviewFlow where (N=16) participants — with little to no experience as reviewers — wrote reviews for two short papers in a counterbalanced manner: one using the ReviewFlow with full scaffolding features and one using the baseline interface that only has the traditional forms of guidance(e.g., review rubrics and an example review). We found that novice reviewers wrote significantly more structured and more comprehensive reviews in the ReviewFlow system than in the Baseline system, evaluated by experts. Novice reviewers wrote slightly more constructive reviews in the ReviewFlow system, but the difference was not significant. Reviewers called out more weaknesses in the paper using ReviewFlow, but they still struggled to provide constructive and actionable suggestions for the authors to overcome the weaknesses.
@@ 2.1. Practices and challenges related to academic peer reviewing
Peer review is an essential step of academic research. As an important step for ensuring the scientific quality of work, peer review has been adopted by most journals and conferences (Price and Flach, 2017). In a typical conference review process, each reviewer needs to evaluate the paper’s quality, make an acceptance decision and provide reviews for their assigned papers (Shah et al., 2018; Shah, 2022). The reviewers are usually experts in the area who have fruitful experience and knowledge to assess or evaluate the quality and contribution of the paper. After reviewers submit the review, a discussion takes place between reviewers and a meta-reviewer, who will carry out the final decision on the acceptance of the paper. Note that conferences may have some differences in their peer-review process.
The number of papers in research communities has increased exponentially in recent years (Wang et al., 2020). While this may be positively viewed as an acceleration of scientific progress, the disparity between growth rates of the submission and reviewer pools also creates more burden for reviewers (McCook, 2006; Stelmakh et al., 2019, 2021a; Shah, 2022). To avoid overloading reviewers, conferences need to find new sources of reviewers as there are not enough experienced reviewers to review all papers (Stelmakh et al., 2021a). These novice and junior reviewers constitute a large fraction of the reviewer pool in computer science conferences (Stelmakh et al., 2021b). For example, using data on the structure of the reviewer pool of the ICML 2020 conference, researchers estimated that approximately 35% of the reviewers were novice reviewers who are junior individuals who self-nominated and satisfied the screening requirements of having one or two papers published in some top ML venues (Stelmakh et al., 2021b). Given this large fraction, conferences need to ensure that newly added junior reviewers do not compromise the quality of the process, that is, are able to write reviews of quality comparable to the experienced reviewers.
@@
Reviewing is a complex, time-consuming, and mentally demanding task (Wang et al., 2020; Shah, 2022). A constructive and comprehensive review can improve the quality of the paper, while a bad, random, dismissive, or biased review brings frustration and anger to reviewers (Wang et al., 2020). To provide a high-quality review, reviewers need to go through a multi-step workflow: first, they need to read and fully understand the contribution of the paper; second, they need to use the domain knowledge to assess the merit of scientific contributions and make a decision on the submission; last, they need to provide an evaluation that covers multiple aspects, including the originality, clarity, validity, etc (Smith, 2006; Jefferson et al., 2002; Xiong and Litman, 2011; Langford and Guzdial, 2015; Hua et al., 2019; Cheng et al., 2020; Gao et al., 2019; Wang et al., 2020; Yuan et al., 2021). To make a fair judgment, reviewers need to have enough background knowledge to grasp the main idea of the paper and evaluate its contribution. More importantly, reviewers need to equip critical thinking skills to think deeply about the author’s judgments, like whether the claims are reasonable and why the approaches are chosen (Facione, 1990; Moore, 2013). A typical peer review not only contains the paper summary and its contribution but also raises weaknesses from different aspects together with constructive feedback or thought-provoking questions. During this process, it requires readers to actively analyze, synthesize, and evaluate the paper content (Moore, 2013).
@@
Researchers are mainly trained to do research while there is little training for the peer review process. Hence, it becomes even more challenging for inexperienced junior reviewers to gain expertise quickly (Shah, 2022). Existing research explored instructional methods to “teach” or “train” junior reviewers. A previous study provided a training video and found that it increased the inter-reviewer agreement, alignment with the scoring rubric, and the amount of time reading the review criteria (Sattler et al., 2015). Another study offers novice reviewers a more guided introduction to the different stages of the reviewing process, such as how to lead a discussion among reviewers, with the goal of helping novices to write better reviews. The results showed that with this guidance on the reviewing stages, novice reviewers can deliver more “above expectation” reviews (Stelmakh et al., 2021a). However, their guidance is only limited to introducing the different parts of the reviewing process, such as rebuttal and discussion, and providing novices opportunities to ask expert questions on the general process, for instance, how to initiate a discussion among reviewers (Stelmakh et al., 2021a). Outside of the general review process, academic peer review is a complex activity that involves both understanding a paper submission’s stated contributions and evaluating whether the paper crosses an acceptable threshold for the research community. To explore how we might support the peer review process, our research starts with a formative study to understand how experts approach this task and what novices see as key challenges.
@@
Previous research has used computational methods to provide support to streamline several parts of the peer review process, such as matching submissions with appropriate peer reviewers, authoring more comprehensive and decisive reviews, as well as review quality assessment (Bartoli et al., 2016; Wang et al., 2020; Hua et al., 2019; Yuan et al., 2021; Arous et al., 2021; Shah et al., 2018). However, there is less empirical study that attempts to scaffold the entire workflow for reviewing academic papers, which includes reading, note-taking, evaluating, decision making, and synthesizing this into a written review.
@@ 2.2. Scaffolding strategies for complex cognitive tasks
To help novices improve problem-solving skills in complex cognitive tasks, cognitive apprenticeship introduces several strategies, including modeling, coaching, scaffolding, and reflection (Vygotsky and Cole, 1978; Collins, 2006). Scaffolding is instructional support provided by experts to promote learning, especially when concepts and skills are being first introduced to novice students (Reiser, 2004; Collins, 2006). These supports include advanced organizers, modeling, worked examples, concept maps, explanations, handouts and prompts (Oakes et al., 2018; Nesbit and Adesope, 2013; Atkinson et al., 2000; Charney and Carlson, 1995; Bui and McDaniel, 2015; Collins, 2006). Previous research shows that effective scaffolding can help novices perform work nearly as well as experts (Yuan et al., 2016; Hui and Sprouse, 2023; Hui et al., 2018). When scaffolding is mediated by technology, including AI-based methods, it creates more opportunities for instructors and learners but also brings more challenges in making the scaffolding contextualized, adaptive, and effective (Hannafin et al., 1999).
@@
Researchers used prompts and guided questions to scaffold learners in the paper reading process (Peng et al., 2022; Yuan et al., 2023; Chen et al., 2022). To facilitate critical paper reading, researchers developed CReBot which interactively asks section-level critical thinking questions for routine paper readers and novice readers. Results showed that the interactive question prompts CReBot provided might not be better than static guidelines for beginners to read. Similarly, researchers developed CriTrainer which can adaptively provide questions in the reading process together with hints and feedback to help readers critically think and comprehend the paper content. Interestingly, on the opposite of CReBot, their result showed that CriTrainer can improve learners’ ability to raise understandable, relevant, and critical questions after the training sessions. Their results highlighted the benefits of its text-specific critical thinking questions provided by the system. However, guided questions used in CReBot and CriTrainer are both template-based, which ignored the contextual information provided by the paper.
@@
Existing research also uses existing examples and templates to scaffold complex writing process (Yuan et al., 2016; Hui and Sprouse, 2023; Hui et al., 2018). In the writing scenario, scholars found that scaffolding can help students learn about form and organization from analyzing the examples and templates (Doan, 2021; Collins, 2006). In the context of writing introductory help requests, providing high-quality examples and expert informed templates can increase learning and writing quality (Hui et al., 2018). Another writing support system used scaffolded annotation and examples as an instruction approach to professional writing students improve the quality of their early-stage drafting in a new genre (Hui and Sprouse, 2023).
In traditional instruction scenarios, experts take a huge amount of time to curate examples, create guidance, or author rubrics (Reiser, 2004). However, experts who created examples still faced the challenges of effectively adapting and contextualizing into the current learning step. Thus, this brings AI and machine intelligence opportunities to provide efficient and context-specific scaffolding in the learning process (Wambsganss et al., 2020, 2021). Previous research has developed multiple methods for using machine intelligence to scaffold complex tasks, including writing and researching (Norkute et al., 2021; Rachatasumrit et al., 2021; Yuan et al., 2022; Gero et al., 2021). For example, researchers developed ArgueTutor, a conversational agent that tutors students with adaptive argumentation feedback in their learning journey using Natural Language Processing (Wambsganss et al., 2021). However, as far as we know, none of the studies that used scaffolding strategies focused on the context of conference peer review writing. Our research explores how we might guide a complex, multi-faceted workflow like peer reviewing and the potential role of AI in creating contextually adaptive scaffolds.
2.3. Leveraging AI to support writing

### 26 Feb 2024 v2

@@ Abstract.
Peer review is a cornerstone of science. Research communities conduct peer reviews to assess contributions and to improve the overall quality of science work. Every year, new community members are recruited as peer reviewers for the first time. How could technology help novices adhere to their community’s practices and standards for peer reviewing? To better understand peer review practices and challenges, we conducted a formative study with 10 novices and 10 experts. We found that many experts adopt a workflow of annotating, note-taking, and synthesizing notes into well-justified reviews that align with community standards. Novices lack timely guidance on how to read and assess submissions and how to structure paper reviews. To support the peer review process, we developed ReviewFlow – an AI-driven workflow that scaffolds novices with contextual reflections to critique and annotate submissions, in-situ knowledge support to assess novelty, and notes-to-outline synthesis to help align peer reviews with community expectations. In a within-subjects experiment, 16 inexperienced reviewers wrote reviews in two conditions: using ReviewFlow and using a baseline environment with minimal guidance. With ReviewFlow, participants produced more comprehensive reviews, identifying more pros and cons. However, they still struggled to provide actionable suggestions to address the weaknesses. While participants appreciated the streamlined process support from ReviewFlow, they also expressed concerns about using AI as part of the scientific review process. We discuss the implications of using AI to scaffold the peer review process on scientific work and beyond.
@@
journalyear: 2024copyright: acmlicensedconference: 29th International Conference on Intelligent User Interfaces; March 18–21, 2024; Greenville, SC, USAbooktitle: 29th International Conference on Intelligent User Interfaces (IUI ’24), March 18–21, 2024, Greenville, SC, USAdoi: 10.1145/3640543.3645159isbn: 979-8-4007-0508-3/24/03ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Peer review is a cornerstone of academic research, ensuring the quality, credibility, and reliability of scientific research (Price and Flach, 2017). The peer review process seeks to assess whether submissions contribute new knowledge to a research community and generate feedback that helps authors improve the quality of their work  (Jefferson et al., 2002; Shah, 2022). Many communities are seeing a rapid increase in submissions (McCook, 2006; Stelmakh et al., 2019; Shah, 2022; Arous et al., 2021; Stelmakh et al., 2021a); while this could be seen as an indicator of scientific progress, it also has increased the pressure on reviewers. To meet increased demand, many research communities recruit a significant number of first-time reviewers or ACs (Associate Chairs) for each review cycle. For example, in 2023, the ACM CHI Conference on Human Factors in Computing Systems reported that more than 50% of ACs were first-time ACs (chi, 2015).
@@
Peer reviewing is a complicated and challenging task. Reviewers need to understand the paper, evaluate the scientific content using domain knowledge, make a fair decision, and compose a comprehensive review to communicate their assessments and recommendations (Shah, 2022). It is a task that requires critical thinking, a deep understanding of the subject, and the ability to provide constructive feedback. Conferences and journals need to ensure that first-time reviewers meet the standards of the research community.
@@
Figure 1. ReviewFlow interface on an example paper. Users can (A) receive section-level contextual cues guided by community criteria; (B) request phrase-level contextual cues adapted to highlight paper content; (C) click the citation to get an in-situ summarization; (D) check the recommended citations not currently cited by the paper; and (E) click to summarize the notes into a high-level outline or expand into a detailed outline
@@
One approach to preparing novices for any complex task is to provide scaffolding, a strategy used in educational settings to support learning and mastery (Reiser, 2004; Collins, 2006). Scaffolding takes many forms, including examples of prior work, templates, or hints to help novices think and potentially perform on par with experts (Holton and Clarke, 2006; Saye and Brush, 2002). Previous studies found that scaffolded examples and templates can even help learners perform similarly to experts in terms of feedback quality (Yuan et al., 2016). For instance, the LetterSmith project developed an approach to aid writing called “scaffolded annotation” which provided key components of the writing tasks and annotated expert examples, and helped professional writing students improve the quality of their early-stage drafts  (Hui and Sprouse, 2023).
Advances in AI and large-language models (LLMs) in particular, have the potential to make scaffolding even more effective because they can detect and adapt to the user’s work context (Rachatasumrit et al., 2022). For example, the CReBot project leverages LLMs to generate and place questions within an academic paper to help induce critical thinking while reading (Peng et al., 2022); results showed that CReBot helped novices read and comprehend the paper content. While LLMs show great potential, they are also known to hallucinate (Bang et al., 2023) and perpetuate systemic biases (Huang et al., 2019; Jakesch et al., 2023) that could negatively impact the effectiveness of scaffolding. While intelligent scaffolding has proven valuable for well-scoped tasks, academic peer review involves reading submissions with a critical eye, synthesizing knowledge, and making a well-justified judgment for acceptance or rejection. Our research investigates how intelligent scaffolds can guide a complex, multi-faceted workflow for academic peer reviewing without biasing the ultimate decisions. This paper explores two key research questions: What are the challenges faced by novices and the strategies adopted by experts during the peer review process? How can we intelligently scaffold the peer review process?
@@
Based on these insights, we developed a prototype – “ReviewFlow” [^1] ReviewFlow Code Repository: https://github.com/LusunHCI/ReviewFlow.git, a platform for writing peer reviews that incorporates intelligent scaffolding to support a workflow for inexperienced reviewers. ReviewFlow incorporates a range of features to facilitate the review process: (1) Contextual cues are embedded questions that help reviewers reflect on the paper. ReviewFlow includes section-level cues guided by the community’s review criteria as well as phrase-level cues that adapt to the paper’s content. (2) In-situ citation recommendations show relevant but non-cited papers that may help reviews assess the novelty compared to existing work. (3) Notes-to-outline synthesis guides reviewers to organize notes and structure reviews to align with community standards. ReviewFlow gathers all the notes left by the reviewer and leverages an LLM to summarize notes into a high-level outline. Reviewers can revise and add detail to the outline while adhering to community standards.
We conducted a within-subjects study to evaluate ReviewFlow where (N=16) participants — with little to no experience as reviewers — wrote reviews for two short papers in a counterbalanced manner: one using the ReviewFlow with all scaffolding features and one using the baseline interface with only traditional forms of guidance (e.g., review rubrics and an example review). We found that novice reviewers wrote significantly more structured and more comprehensive reviews in the ReviewFlow system than in the Baseline system, as evaluated by experts. Novice reviewers wrote slightly more constructive reviews in the ReviewFlow system, but the difference was not significant. Reviewers called out more weaknesses in the paper using ReviewFlow, but they still struggled to provide actionable suggestions for the authors to address the weaknesses.
@@ 2.1. Practices and challenges related to academic peer reviewing
As an important step for ensuring the scientific quality of work, peer review has been adopted by most journals and conferences (Price and Flach, 2017). In a typical conference review process, each reviewer needs to evaluate the paper’s quality, make an acceptance decision and provide reviews for their assigned papers (Shah et al., 2018; Shah, 2022). The reviewers are usually experts in the area who have fruitful experience and knowledge to assess or evaluate the quality and contribution of the paper. After reviewers submit the review, a discussion takes place between reviewers and a meta-reviewer, who will carry out the final decision on the acceptance of the paper.
The number of papers in research communities has increased exponentially in recent years (Wang et al., 2020). While this may be positively viewed as an acceleration of scientific progress, the disparity between growth rates of the submission and reviewer pools also creates more burden for reviewers (McCook, 2006; Stelmakh et al., 2019; Stelmakh et al., 2021a; Shah, 2022). To avoid overloading reviewers, conferences need to find new sources of reviewers as there are not enough experienced reviewers to review all papers (Stelmakh et al., 2021a). These novice and junior reviewers constitute a large fraction of the reviewer pool in computer science conferences (Stelmakh et al., 2021b). For example, in 2023, the ACM CHI Conference on Human Factors in Computing Systems reported that more than 50% of ACs who reviewed papers are first-time ACs (chi, 2015). Given this large fraction, conferences need to ensure that newly added junior reviewers do not compromise the quality of the process, that is, are able to write reviews of quality comparable to the experienced reviewers.
@@
Reviewing is a time-consuming and mentally demanding task (Wang et al., 2020; Shah, 2022). A constructive and comprehensive review can improve the quality of the paper, while a bad, random, dismissive, or biased review brings frustration and anger to authors (Wang et al., 2020). To provide a high-quality review, reviewers need to go through a multi-step workflow, including understanding the contribution of the paper, accessing the merit of scientific contribution and providing an evaluation together with a comprehensive written review (Smith, 2006; Jefferson et al., 2002; Xiong and Litman, 2011; Langford and Guzdial, 2015; Hua et al., 2019; Cheng et al., 2020; Gao et al., 2019; Wang et al., 2020; Yuan et al., 2021). To make a fair judgment, reviewers need to have enough background knowledge to grasp the main idea of the paper and evaluate its contribution. More importantly, reviewers need to equip critical thinking skills to think deeply about the author’s judgments, like whether the claims are reasonable and why the approaches are chosen (Facione, 1990; Moore, 2013). A typical peer review not only contains the paper summary and its contribution but also raises weaknesses from different aspects together with constructive feedback or thought-provoking questions. During this process, it requires readers to actively analyze, synthesize, and evaluate the paper content (Moore, 2013).
@@
Researchers are typically trained extensively in conducting the research itself, but they often lack formal instructions in the peer review process. Hence, it becomes even more challenging for inexperienced junior reviewers to gain expertise quickly (Shah, 2022). Existing research explored instructional methods to “teach” or “train” junior reviewers. A previous study provided a training video and found that it increased the inter-reviewer agreement, alignment with the scoring rubric, and the amount of time reading the review criteria (Sattler et al., 2015). Another study offers novice reviewers a more guided introduction to the different stages of the reviewing process, such as how to lead a discussion among reviewers, to help novices write better reviews. The results showed that with this guidance on the reviewing stages, novice reviewers could deliver more “above expectation” reviews (Stelmakh et al., 2021a). However, their guidance is only limited to introducing the different parts of the reviewing process, such as rebuttal and discussion, and providing novices opportunities to ask expert questions on the general process (Stelmakh et al., 2021a). Outside of the general review process, academic peer review is a complex activity that involves both understanding a paper submission’s stated contributions and evaluating whether the paper crosses an acceptable threshold for the research community. To explore how we might support the peer review process, we start with a formative study to understand how experts approach this task and what novices see as key challenges.
@@
Previous research has used computational methods to provide support to streamline several parts of the peer review process, such as matching submissions with appropriate reviewers or assessing review quality (Bartoli et al., 2016; Wang et al., 2020; Hua et al., 2019; Yuan et al., 2021; Arous et al., 2021; Shah et al., 2018). However, fewer empirical studies that attempt to scaffold the entire workflow for reviewing academic papers, which includes reading, note-taking, evaluating, decision making, and synthesizing this into a written review.
@@ 2.2. Scaffolding strategies for complex cognitive tasks
To help novices improve problem-solving skills in complex cognitive tasks, cognitive apprenticeship introduces several strategies, including modeling, coaching, scaffolding, and reflection (Vygotsky and Cole, 1978; Collins, 2006). Scaffolding is instructional support provided by experts to promote learning, especially when concepts and skills are being first introduced to novice students (Reiser, 2004; Collins, 2006). These supports include advanced organizers, modeling, worked examples, concept maps, explanations, handouts, and prompts (Oakes et al., 2018; Nesbit and Adesope, 2013; Atkinson et al., 2000; Charney and Carlson, 1995; Bui and McDaniel, 2015; Collins, 2006). Previous research shows that effective scaffolding can help novices perform work nearly as well as experts (Yuan et al., 2016; Hui and Sprouse, 2023; Hui et al., 2018). When scaffolding is mediated by technology, including AI-based methods, it creates more opportunities for instructors and learners but also brings more challenges in making the scaffolding contextualized, adaptive, and effective (Hannafin et al., 1999).
@@
Researchers used prompts and guided questions to scaffold learners in the paper reading process (Peng et al., 2022; Yuan et al., 2023; Chen et al., 2022). To facilitate critical paper reading, researchers developed CReBot which interactively asks section-level critical thinking questions for routine paper readers and novice readers. Results showed that the interactive question prompts CReBot provided might not be better than static guidelines for beginners to conduct critical thinking. Furthermore, researchers developed CriTrainer which can adaptively provide questions in the reading process together with hints and feedback to help readers critically think and comprehend the paper content. Interestingly, on the opposite of CReBot, their result showed that CriTrainer can improve learners’ ability to raise understandable, relevant, and critical questions after the training sessions. Their results highlighted the benefits of its text-specific critical thinking questions provided by the system. However, guided questions used in CReBot and CriTrainer are both template-based and did not fully use the user-selected content on the paper.
@@
Existing research also used existing examples to scaffold complex writing processes (Yuan et al., 2016; Hui and Sprouse, 2023; Hui et al., 2018). Scholars found that scaffolding can help students learn about form and organization by analyzing the examples and templates (Doan, 2021; Collins, 2006). In the context of writing introductory help requests, providing high-quality examples and expert-informed templates can increase learning and writing quality (Hui et al., 2018). Another writing support system used “scaffolded annotation” that broke down examples into each component to help professional writing students improve their early-stage drafts (Hui and Sprouse, 2023). In traditional instruction scenarios, experts take a large amount of time to curate examples, create guidance, or author rubrics (Reiser, 2004). However, experts who created examples still faced the challenges of effectively adapting and contextualizing into the current learning step.
2.3. Leveraging AI scaffolding to support writing